# Let’s Get to the Point: LLM-Supported Planning, Drafting, and Revising of Research-Paper Blog Posts

### v1

@@ Abstract.
Research-paper blog posts help scientists disseminate their work to a larger audience, but translating papers into this format requires substantial additional effort. Blog post creation is not simply transforming a long-form article into a short output, as studied in most prior work on human-AI summarization. In contrast, blog posts are typically full-length articles that require a combination of strategic planning grounded in the source document, well-organized drafting, and thoughtful revisions. Can tools powered by large language models (LLMs) assist scientists in writing research-paper blog posts? To investigate this question, we conducted a formative study (N=6) to understand the main challenges of writing such blog posts with an LLM: high interaction costs for 1) reviewing and utilizing the paper content and 2) recurrent sub-tasks of generating and modifying the long-form output. To address these challenges, we developed Papers-to-Posts, an LLM-powered tool that implements a new Plan-Draft-Revise workflow, which 1) leverages an LLM to generate bullet points from the full paper to help users find and select content to include (Plan) and 2) provides default yet customizable LLM instructions for generating and modifying text (Draft, Revise). Through a within-subjects lab study (N=20) and between-subjects deployment study (N=37 blog posts, 26 participants) in which participants wrote blog posts about their papers, we compared Papers-to-Posts to a strong baseline tool that provides an LLM-generated draft and access to free-form LLM prompting. Results show that Papers-to-Posts helped researchers to 1) write significantly more satisfying blog posts and make significantly more changes to their blog posts in a fixed amount of time without a significant change in cognitive load (lab) and 2) make more changes to their blog posts for a fixed number of writing actions (deployment).
@@ 1. Introduction
Sharing scientific ideas outside of research papers is important for increasing awareness and impact (Jucan and Jucan, 2014; Burns et al., 2003). Scientists often seek different ways to communicate their work to broader audiences, such as through tweets or blog posts (Gero et al., 2021; Long et al., 2023; Williams et al., 2022; Holmberg and Thelwall, 2014); indeed, tweeting about research papers has been shown to raise their citation count (Luc et al., 2021). While some tools have been created to help author short-form derivatives like tweets (Long et al., 2023), significant challenges remain for scientists seeking to create longer-form derivatives like blog posts, which require more extensive planning, drafting, and revising writing processes (Flower and Hayes, 1981). In the planning step, the blog post author must review the paper, itself a long document, to find and select several pieces of information to include in the blog post. In the drafting and revising steps, there are many sub-tasks for completing the full blog post, such as creating an engaging hook or reporting results with sufficient detail but in simpler terms than the paper.
@@
Integrating large language models (LLMs) in writing tools offers a promising avenue for improving writing outcomes; however, existing approaches are not designed to help create long-form derivatives like blog posts from long documents like academic papers. Prior work in human-LLM text summarization has focused on supporting short-form summaries, a sentence or paragraph in length (Zhang et al., 2023b; Lai et al., 2022). In the broader area of human-LLM writing, prior work has looked into providing assistance for writing long-form creative texts such as stories (Chung et al., 2022; Goldfarb-Tarrant et al., 2019; Singh et al., 2022) and screenplays (Mirowski et al., 2023), but these tasks do not take a long-form text input. Prior work in human-LLM scientific writing has investigated how LLM tools can assist scientists in writing short-form texts such as tweetorials (Long et al., 2023) and extended metaphors (Kim et al., 2023), but to the best of our knowledge, no work has sought to develop an LLM tool for creating long-form scientific texts, such as research-paper blog posts.
Still, prior work has shown that LLM tools can support planning, drafting, and revising for different forms of writing, such as creative stories (Goldfarb-Tarrant et al., 2019), short summaries (Zhang et al., 2023b), and short scientific texts (Long et al., 2023). Thus, in this paper we ask: how best can an LLM tool provide assistance in dealing with the challenges of writing long-form summaries of scientific long documents? We designed a technology probe of a simple writing interface with preliminary LLM-based features to better understand the costs and benefits of interacting with LLMs for the purpose of planning, drafting, and revising a research-paper blog post. The technology probe provides the paper text, an LLM-generated bullet point outline for the paper, a drafting space to write the blog post in sections, and access to an LLM prompting interface for writing tasks. Through a formative study (N=6), we gathered feedback on how users wish to leverage LLMs in this context and common interaction challenges. For one, the participants valued having a paper outline for finding and selecting relevant information from the paper to include in the blog post, but they struggled with understanding and utilizing it (e.g., understanding how the bullet points were connected to the paper text). Also, participants desired to provide the LLM recurring instructions for generating and modifying text that were not well-supported by the tool (e.g., rewrite a paragraph to better capture readers’ attention, continue writing from a particular sentence).
@@
Based on the identified challenges, we iterated on the technology probe to develop the prototype tool Papers-to-Posts for LLM-assisted paper blog post writing. The tool follows our Plan-Draft-Revise workflow (Figure 1). In the Plan phase, the user may control text generation by selecting a subset of LLM-generated bullet points corresponding to key points in the source; Papers-to-Posts is the first LLM-powered long-document summarization tool to support direct control of content selection through a source outline, maintaining the source’s narrative structure. In the Draft phase, the user may supply additional instructions and preview LLM generations that use the selected bullet points, instructions, and blog post text thus far. In the Revise phase, they may use an LLM to further refine text with macros that encapsulate frequent editing transformations (e.g., to have less jargon, to be more dramatic).
@@
We evaluated Papers-to-Posts through two studies in which participants wrote blog posts for their own papers, and we compared our tool to a baseline tool that provided an LLM-generated draft and integrated prompting. In a within-subjects lab study (N=20) in which researchers wrote blog posts for two of their own papers, one per tool, we found that, by helping authors to plan, draft, and revise their blog post, Papers-to-Posts led to significantly higher writer satisfaction with the blog post output alongside significantly greater change to the provided draft, without increasing task completion time or cognitive load. To study Papers-to-Posts in a more realistic setting, we also conducted an unmonitored, between-subjects deployment study (N=37), in which participants using Papers-to-Posts again made greater change to the provided drafts for a given amount of time spent or number of writing actions. Furthermore, through quantitative survey ratings and behavioral analysis as well as qualitative short answers and semi-structured interviews across the two studies, we observe that an LLM-generated paper outline and affordances for streamlining common long-form text generation and modification sub-tasks contributed to this increased satisfaction. In discussing our results, we also reflect on design implications for other human-LLM scientific long-document summarization tools with long-form output.
In summary, we make the following main contributions:
A formative study (N=6) producing two primary design goals for an LLM-powered tool for writing research-paper blog posts: (1) help users review and select content from a long source document during the planning stage, and (2) transparently streamline the process of instructing the LLM to generate and modify text during the drafting and revising stages.
Papers-to-Posts, a prototype tool for writing research-paper blog posts, which addresses the design goals identified in the formative study and implements the Plan-Draft-Revise workflow (Figure 1).
A within-subjects lab study (N=20 participants) providing evidence that Papers-to-Posts improves upon a powerful baseline (LLM draft and free-form prompting). Papers-to-Posts led to significantly higher writer satisfaction with the final blog post (with comparable task completion time and cognitive load) alongside significantly greater changes to the initial draft.
A field deployment study (N=37 blog posts, 26 participants) comparing Papers-to-Posts to the same strong baseline show that the same benefits hold in realistic scenarios where participants were writing blog posts on their own and sharing them publicly at the end of the study.
Design implications for other human-LLM scientific long-document summarization tools with long-form output: avoid solely post-editing interactions, present long-document outlines to support content selection, and include transparent affordances to streamline sub-tasks in generating and modifying long-form output text.
@@ 2. Related Work
2.1. Human-LLM Summarization
Although there is much work on automatic text summarization (Koh et al., 2022; El-Kassas et al., 2021), issues remain with the output summaries, such as inaccuracies and style misalignment (Lu et al., 2023; Koh et al., 2022). To address these issues, other works have looked into having a human and LLM work together to generate a summary (Lai et al., 2022; Moramarco et al., 2021). According to Cheng et al., there are five common methods of human-AI summarization: guiding AI output, selecting or rating AI output, post-editing of AI-generated summaries, interactive editing, and writing with AI assistance (Cheng et al., 2022). Our tool Papers-to-Posts provides affordances related to guiding and post-editing AI output.
Prior work in human-AI summarization has mainly focused on direct model guidance (e.g., utilizing concepts/queries (Zhang et al., 2023b; Bayatmakou et al., 2022; Avinesh and Meyer, 2017; Avinesh et al., 2018; Ding et al., 2023; Wang et al., 2023b), typing (Xie et al., 2023), ratings (Gao et al., 2020; Bohn and Ling, 2021), gaze (Bohn and Ling, 2021)) or simply post-editing (e.g., (Lai et al., 2022; Moramarco et al., 2021)). A few works have supported summarization based on user-selected points from the source (Bayatmakou et al., 2022; Yan et al., 2011; Zhang et al., 2011; Wang et al., 2023a, b), but the points are ranked or clustered by the system without the same narrative structure as the source document. Another work proposed but did not implement or evaluate model guidance through in-context content selection (Gehrmann et al., 2019). Our approach allows model guidance directly grounded in the input source while maintaining the source narrative structure.
Providing more direct control of input content is useful when the writer has significant knowledge of the source and can use reminders of its contents to select what to include in the summary. We focus on summarizing one’s own content, unlike most prior work, with the exception of work on summarizing medical notes that used the post-editing technique (Moramarco et al., 2021; Cai et al., 2022). Thus, we take advantage of the writer’s knowledge of the source document by having them select content in context of the source’s layout.
@@ 2.1. Human-LLM Summarization
There are several works on human-LLM summarization of long documents (Zhang et al., 2011; Gehrmann et al., 2019) and some for scientific long documents in particular (Zarinbal et al., 2019; Bayatmakou et al., 2022; Zhang et al., 2023b). Long documents provide a unique challenge for human-LLM summarization, as there is a great deal of information for the human to digest. As writing research-paper blog posts entails summarizing portions of a research paper, this work seeks to support human-LLM long-document summarization.
2.2. Human-LLM Writing
While LLMs have impressive capabilities with respect to quickly generating large amounts of text, interacting with swaths of text can be difficult to manage without additional affordances. Popular chatbot LLMs such as chatGPT (OpenAI, [n. d.]) and BARD (Google, [n. d.]) offer some options for the user to guide their output beyond general prompting. Both provide example inputs for the user to select. chatGPT includes the option to provide custom instructions for the questions ”What would you like chatGPT to know about you to provide better responses?” and ”How would you like chatGPT to respond?”. BARD allows the user to modify its generated text to be shorter, longer, simpler, more casual, more professional, or modified based on a custom prompt. Tools from prior research also provide users with means to guide LLM writing beyond general prompting, such as keyword specifications (Zhang et al., 2023b), style specifications (Ghazvininejad et al., 2017; Coenen et al., 2021), and input modification (Zhang et al., 2011).
@@ 2.2. Human-LLM Writing
With respect to planning text, a unique aspect of our work is providing an LLM-generated outline for source content selection. There are a couple works in human-LLM writing that have utilized outlines, but they focused on outlining output content rather than input content. Zhang et al. introduce the tool VISAR, a human-AI argumentative writing assistant that allows users to iterate on their draft with adjustments to a visual outline of the draft (Zhang et al., 2023a). Dang et al. introduce a text editor that generates summaries of each paragraph written in the editor to comprise a reverse outline of the writing (Dang et al., 2022). The summaries then become affordances to help writers revise their writing (e.g., restructuring the writing). Our work also generates a reverse outline but as an affordance to guide summary generation based on the input content. For summarizing a long source document, an outline is particularly useful, as reviewing the entire source may require substantial time and effort.
With respect to revising text, like BARD (Google, [n. d.]) and other LLM-powered writing tools (NotionAI, [n. d.]), our final tool allows writers to utilize preset modifications, but some of these are set to be more specific to blog post writing. Closely related to our work, Laban et al. present an editing interface to help users revise their writing with LLM edit suggestions that can be integrated into their writing (Laban et al., 2023). This editing interface focuses solely on the revising step and on smaller edits than our tool.
@@ 2.3. Human-LLM Scientific Writing
Several works have investigated how to support human-LLM writing. Most of these works focus on tools for creative writing (Singh et al., 2022; Chung et al., 2022; Mirowski et al., 2023), but a few cater to scientific writing. A few works have investigated human-LLM scientific writing for broader science communication tasks that do not involve a source document. Gero et al. present a system for generating scientific ”sparks”, or inspiring sentences (Gero et al., 2022). Long et al. examine how LLM scaffolding can help people generate relatable hooks for complex scientific topics (Long et al., 2023), and Kim et al. investigate how science writers can generate extended metaphors for scientific ideas with the help of an LLM (Kim et al., 2023). The couple works that have explored human-AI scientific summarization do not allow for in-context selection of source content but rely on selection of concepts (Zhang et al., 2023b) or out-of-context sentences extracted based on queries (Bayatmakou et al., 2022).
These previously studied forms of human-LLM scientific writing are the length of a sentence or paragraph. Based on prior work (Mehlenbacher, 2019), popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]; Medium, [n. d.]), and feedback from three communications experts, we determined a set of guidelines for writing research-paper blog posts, which indicate that the shortest blog posts are recommended to be around 500 words in length (Medium, [n. d.]; Kudos, [n. d.]; Patrick Dunleavy, [n. d.]), making blog posts a long-form output. Because research-paper blog posts are longer than a sentence or a paragraph, they require coordinating more writing sub-tasks, such as rewriting a sentence about a method to be easier for a broader audience to understand or summarizing a paper’s results section.

### v2

@@ Abstract.
Research-paper blog posts help scientists to disseminate their work to a larger audience, but translating scientific long documents into long-form summaries like blog posts raises unique challenges: 1) planning what paper content to include in the blog post, 2) drafting the selected content in sections amenable to a paper blog post, and 3) revising the blog post to be scientifically accurate but also concise, easy to understand, and engaging. Can we harness the power of large language models (LLMs) to assist researchers with these challenges? To investigate this question, we developed Papers-to-Posts, an LLM-powered tool that implements a new Plan-Draft-Revise workflow for mixed-initiative long-form paper summarization. An LLM-generated paper outline with pre-selected yet adjustable bullet points helps users to plan what information to include. Meanwhile, customizable LLM instructions support drafting the text with a suitable structure and revising the text to have an appropriate tone. Through two studies, we compared Papers-to-Posts to a strong baseline tool that provides an LLM-generated draft and access to free-form LLM prompting, and we found that Papers-to-Posts improved researchers’ editing power. In a within-subjects lab study (N=20 participants), Papers-to-Posts led participants to make significantly more change to initial LLM drafts within a fixed amount of time and to be significantly more satisfied with their final blog post, without increasing cognitive load. Furthermore, in a between-subjects deployment study (N=37 blog posts, 26 participants), Papers-to-Posts led participants to make more change to initial LLM drafts within a given amount of time as well as writing actions, without decreasing satisfaction with the final blog posts or increasing cognitive load.
@@ 1. Introduction
Sharing scientific ideas outside of research papers is important for increasing awareness and impact (Jucan and Jucan, 2014; Burns et al., 2003). Scientists often seek different ways to communicate their work to broader audiences, such as through tweets or blog posts (Gero et al., 2021; Long et al., 2023; Williams et al., 2022; Holmberg and Thelwall, 2014); indeed, tweeting about research papers has been shown to raise their citation count (Luc et al., 2021). While some tools have been created to help author short-form derivatives like tweets (Long et al., 2023), significant challenges remain for scientists seeking to create longer-form derivatives like blog posts, which require more extensive planning, drafting, and revising writing processes (Flower and Hayes, 1981). Integrating large language models (LLMs) in writing tools offers a promising avenue for improving editing power (i.e., change in blog post per minute or writing action). However, existing approaches are not designed to help create long-form derivatives like blog posts from scientific long documents like academic papers.
@@
Prior work in human-LLM text summarization (Zhang et al., 2023b; Lai et al., 2022) and human-LLM scientific writing (Long et al., 2023; Kim et al., 2023) has focused on supporting short-form summaries, a sentence or paragraph in length. In the broader area of human-LLM writing, prior work has looked into providing assistance for writing long-form creative texts (Chung et al., 2022; Goldfarb-Tarrant et al., 2019; Singh et al., 2022; Mirowski et al., 2023), but these tasks do not take a long-form text input. To the best of our knowledge, no work has developed and evaluated a mixed-initiative, LLM-powered tool for generating long-form summaries of scientific long documents, such as research-paper blog posts. Shamsabadi and D’Souza presented a prompt-based tool to support different research tasks including writing blog posts; however, the tool’s main function is to generate a personalized prompt for the user to then provide to an LLM, and the tool was not evaluated (Shamsabadi and D’Souza, 2024). Still, prior work has shown that LLM-powered tools can support planning, drafting, and revising different writing forms, such as creative stories (Goldfarb-Tarrant et al., 2019), short summaries (Zhang et al., 2023b), and short scientific texts (Long et al., 2023). Thus, in this paper we ask: can an LLM-powered tool provide assistance in dealing with the challenges of writing long-form summaries of scientific long documents?
As indicated by the many guidelines that discuss the content, formatting, and language of academic blog posts (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]), there are multiple challenges specific to summarizing a research paper in a blog post. The first distinct challenge is related to the input being a long document. The writer must sift through the paper to determine what information to carry over to the blog post, without missing important details. The two other main challenges are related to the output being long-form and for science communication. The writer must organize key points in sections amenable to a scientific blog post. Also, the writer must ensure that their writing is scientifically accurate like the paper but more concise, easy to understand, and eye-catching than the paper.
@@
To help researchers write research-paper blog posts, we developed Papers-to-Posts, a mixed-initiative tool designed based on guidelines for academic blog post writing (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]; Mehlenbacher, 2019) as well as pilot study feedback. The tool follows our Plan-Draft-Revise workflow (Figure 1). In the Plan phase, the user may adjust the tool’s selection of key bullet points within an LLM-generated paper outline. Inspired by prior work on LLM-supported reverse outlining (Dang et al., 2022), Papers-to-Posts is, to the best of our knowledge, the first LLM-powered long-document summarization tool to support direct control of content selection through an LLM-generated document outline, maintaining the source’s narrative structure. In the Draft phase, the user may customize preset instructions (e.g., include a hook, make one paragraph) to the LLM for each blog post section. The LLM then uses the selected bullet points and adjusted instructions to generate text for the given section. In the Revise phase, the user may further refine text with LLM-powered macros that encapsulate frequent editing transformations for scientific blog post writing (e.g., to reduce jargon, to better capture attention).
@@
We evaluated Papers-to-Posts through two studies in which participants wrote blog posts for their own papers. In these studies, we compared our tool to a strong baseline tool–an LLM-generated draft and access to free-form LLM prompting. In a within-subjects lab study (N=20 participants) in which researchers wrote blog posts for two papers (one per tool), we found that Papers-to-Posts led to significantly greater change to the provided draft within a fixed amount of time alongside significantly higher writer satisfaction with the final blog post, without significantly increasing cognitive load. To study Papers-to-Posts in a more realistic setting, we also conducted an unmonitored, between-subjects deployment study (N=37 blog posts, 26 participants). Again, participants using Papers-to-Posts made greater change to the provided drafts for a given amount of time and for a given number of writing actions taken. Through surveys, interviews, and logs from across the two studies, we also observed evidence that an interactive LLM-generated paper outline for planning writing as well as preset but customizable LLM instructions for drafting and revising writing contributed to this increased editing power. In discussing our results, we reflect on implications for other human-LLM tools for research-paper summarization and long-form scientific communication: avoid solely post-editing interactions, provide an interactive outline of the long document for planning the summary, and include preset but adjustable LLM instructions for generating and modifying the writing.
In summary, we make the following contributions:
Papers-to-Posts, a tool for writing research-paper blog posts that implements the Plan-Draft-Revise workflow (Figure 1 and Section 3), including providing an interactive LLM-generated paper outline for planning the blog post, section-specific and customizable LLM instructions for drafting the blog post, and macros for revising text style to suit the blog post.
Findings from a within-subjects lab study (N=20 participants) showing that Papers-to-Posts significantly increases researchers’ editing power (i.e., change in writing within a fixed amount of time) (Figure 5a and Section 5.1.1), while significantly increasing satisfaction with the output blog post (Figure 6a and Section 5.2.1) and without significantly impacting cognitive load.
Findings from a between-subjects deployment study (N=37 blog posts, 26 participants) demonstrating that Papers-to-Posts increases researchers’ editing power (i.e., change in writing within an amount of time or writing actions) (Figure 5b-c and Section 5.1.2), without decreasing satisfaction with the output blog post or increasing cognitive load.
@@ 2. Related Work
2.1. Human-LLM Text Summarization
@@
Although there is much work on automatic text summarization (Koh et al., 2022; El-Kassas et al., 2021), issues remain with the output summaries, such as inaccuracies and style misalignment (Lu et al., 2023; Koh et al., 2022). To address these issues, other works have looked into having a human and LLM work together to generate a summary (Lai et al., 2022; Moramarco et al., 2021). There are several works on human-LLM summarization of long documents (Zhang et al., 2011; Gehrmann et al., 2019) and some for scientific long documents in particular (Zarinbal et al., 2019; Bayatmakou et al., 2022; Zhang et al., 2023b; Cheng et al., 2023). Long documents provide a unique challenge for human-LLM summarization, as there is a great deal of information for the human to digest. As writing research-paper blog posts entails summarizing portions of a research paper, this work seeks to support human-LLM long-document summarization.
According to Cheng et al., there are five common methods of human-AI summarization: guiding AI output, selecting or rating AI output, post-editing AI-generated summaries, interactive editing, and writing with AI assistance (Cheng et al., 2022). Our tool Papers-to-Posts provides affordances related to guiding and post-editing AI output. Prior work in human-AI summarization has mainly focused on model guidance through various means (e.g., concepts/queries (Zhang et al., 2023b; Bayatmakou et al., 2022; Avinesh and Meyer, 2017; Avinesh et al., 2018; Ding et al., 2023; Wang et al., 2023b), typing (Xie et al., 2023), ratings (Gao et al., 2020; Bohn and Ling, 2021), gaze (Bohn and Ling, 2021)) or simply post-editing (e.g., (Lai et al., 2022; Moramarco et al., 2021)). A few works have supported summarization based on user-selected points from the source artifact (Bayatmakou et al., 2022; Yan et al., 2011; Zhang et al., 2011; Wang et al., 2023a, b), but the points available for selection are ranked or clustered by the system without the same narrative structure as the source. Another work proposed but did not implement or evaluate model guidance through in-context content selection (Gehrmann et al., 2019). Our approach allows model guidance directly grounded in the input source while maintaining the source narrative structure.
Providing more direct control of input content is useful when the writer has significant knowledge of the source and can use reminders of its contents when selecting what to include in the summary. We focus on summarizing one’s own content, unlike most prior work; the exception is work on summarizing medical notes, which used the post-editing technique (Moramarco et al., 2021; Cai et al., 2022). We take advantage of the writer’s knowledge of the source document by having them select content within the context of the source’s layout.
2.2. Outlines in Human-LLM Writing
@@
With respect to planning blog post text, a unique aspect of our work is providing an interactive LLM-generated outline for source content selection. There are some works in human-LLM writing that have utilized outlines for writing, but they focused on outlining output content rather than input content. For example, Zhang et al. introduced the tool VISAR, a human-AI argumentative writing assistant that allows users to iterate on their draft with adjustments to a visual outline of the draft (Zhang et al., 2023a), and Lu et al. created a tool to support writing workplace social media posts that includes a feature to help generate an outline for the post (Lu et al., 2024). Most related to our work, Dang et al. introduced a text editor that generates summaries of each paragraph written in the editor to comprise a reverse outline of the writing (Dang et al., 2022). The summaries then become affordances to help writers revise their writing (e.g., restructuring the writing). Our work also generates a reverse outline, but it is an outline of the input paper for users to interactively guide blog post generation. For summarizing a long source document, an outline is particularly useful, as reviewing the entire source may require substantial time and effort.
2.3. Preset, Customizable LLM Instructions in Human-LLM Writing
For drafting blog post text, our tool presents LLM instructions preset to generate four unique research-paper blog post sections; the user may then adjust these section-specific instructions in multiple manners to draft text that better aligns with their needs. Prior works also break text into chunks to support drafting and revising writing. For example, Laban et al. present an editing interface that provides edit suggestions for relatively small chunks of writing (Laban et al., 2023), while Dhillon et al. provide next-sentence and next-paragraph suggestions to continue writing (Dhillon et al., 2024). There are also prior works on human-LLM writing that, like ours, allow users to provide customized instructions to LLMs for drafting text (e.g., (Lim et al., 2024; Coenen et al., 2021)). Our customizable LLM instructions for drafting text are specifically designed to support writing common sections of research-paper blog posts.
For revising blog post text, our tool provides preset macros as well as the ability to perform custom text modifications. Several existing tools (e.g., (Google, [n. d.]; NotionAI, [n. d.]; Lu et al., 2024; Coenen et al., 2021; Reza et al., 2024)) also allow users to modify their text with macros for actions such as adjusting the text’s length, simplicity, or professionalism, and some allow custom modifications. However, our tool’s macros are designed specifically for blog post writing.
@@ 2.4. Human-LLM Scientific Writing
Several works have investigated how to support human-LLM writing. Most of these works focus on tools for creative writing (Singh et al., 2022; Chung et al., 2022; Mirowski et al., 2023), but a few cater to scientific writing. A few works have investigated human-LLM scientific writing for broader science communication tasks that do not involve a source document. Gero et al. present a system for generating scientific “sparks”, or inspiring sentences (Gero et al., 2022). Long et al. examine how LLM scaffolding can help people generate relatable hooks for complex scientific topics (Long et al., 2023), and Kim et al. investigate how science writers can generate extended metaphors for scientific ideas with the help of an LLM (Kim et al., 2023). The few works that have explored human-AI scientific summarization do not allow for in-context selection of source content but rather allow users to select concepts (Zhang et al., 2023b), out-of-context sentences extracted based on queries (Bayatmakou et al., 2022), suggested summary content (Lim et al., 2024). Lim et al.’s Q-and-A research-paper summarization tool AQUA also allows users to prompt the LLM with relevant questions (Lim et al., 2024). Meanwhile, Google’s tool NotebookLM for general human-AI document summarization does not permit any content selection (Google, [n. d.]).
Based on prior work (Mehlenbacher, 2019), popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]), and feedback from three communications experts, we determined a set of guidelines for writing research-paper blog posts, which indicate that the shortest blog posts are recommended to be around 500 words in length, making blog posts a long-form output. Because research-paper blog posts provide long-form scientific communication, drafting them requires LLM instructions unique to common blog post sections, and revising them requires balancing the writing style for accuracy, engagement, ease of understanding, and concision. Most of the previously cited forms of human-LLM scientific writing above are the length of a sentence or paragraph. The exceptions are AQUA’s Q-and-A paper summary (Lim et al., 2024) and NotebookLM’s many summary forms (e.g., FAQ, study guide, briefing doc, podcast). However, AQUA’s scaffolding is limited to suggested questions and answers as well as a space for user-provided additional prompting, and NotebookLM provides no scaffolding to support customization of their document summaries. Furthermore, our work is unique in its dedication to supporting mixed-initiative writing of research-paper blog posts.